OpenStudy (anonymous):

Can someone help me solve for the smaller segment of a circle. chord 3sqr2 and angle is 90 degrees

OpenStudy (phi):

Do you have a picture ?

OpenStudy (phi):

and exactly what is the question ?

OpenStudy (anonymous):

yes,

OpenStudy (phi):

is it like this? 

OpenStudy (anonymous):

yes that's the picture

OpenStudy (phi):

and exactly what is the question ?

OpenStudy (anonymous):

find the area of the smaller segmebt of the cicle. round your answer to two decimal points

OpenStudy (phi):

First, you need to find the radius. Notice you have a right angle so you have a right triangle do you know the pythagorean theorem ?

OpenStudy (anonymous):

c=asq2 + bsqr2

OpenStudy (phi):

yes, now to use that formula we have to label the triangle's sides. we can label its legs "r" (for radius) because both legs are a radius. what is c (hypotenuse) in our triangle ?

OpenStudy (anonymous):

3sqr2

OpenStudy (phi):

ok. we should tweak your version of pythagoras it is \[ c^2 = a^2 + b^2 \] to use it , replace c with \(3 \sqrt{2} \) replace a with r replace b with r can you do that ?

OpenStudy (anonymous):

ok, i did that and equation is 18=rsqr2+rsqr2

OpenStudy (phi):

yes if you use the equation editor it looks nicer \[ 18 = r^2 + r^2 \] if you have 1 r squared and add another r squared how many r^2 do we have ?

OpenStudy (anonymous):

2

OpenStudy (phi):

which means \[ 18 = r^2 + r^2 \] simplifies to \[ 18 = 2r^2 \] can you solve for r ? The next step is divide both sides by 2

OpenStudy (anonymous):

r=3

OpenStudy (phi):

The next step is what is the formula for the area of a circle ?

OpenStudy (anonymous):

\[pirsq\]

OpenStudy (phi):

A = pi r^2 (use ^ to show 2 is an exponent) so the area of the whole circle is A= pi * 3^2 = 9 pi

OpenStudy (anonymous):

ok thank u

OpenStudy (phi):

next, the "sector" is what fraction of the circle ?

OpenStudy (phi):

if the whole circle is a pie, the sector is what fraction of the whole pie ?

OpenStudy (phi):

can you guess ?

OpenStudy (anonymous):

1/4

OpenStudy (phi):